OpenStudy (anonymous):

A wheel makes 2 4/15 revolutions per minute. If it rotates for 70 minutes, how many revolutions does it make?

OpenStudy (anonymous):

158.67

OpenStudy (anonymous):

or 476/3

OpenStudy (radar):

Total revolutions=2 4/15 X 70 = 34/15 X 70= (34 X70)/3 = 158.67

OpenStudy (anonymous):

simplify please & the answer needs to be typed as a whole number,proper fraction, or mixed number thank you.

OpenStudy (radar):

The answer as a mixed number would be 158 2/3. did you follow the steps in the third post up from this post?

OpenStudy (anonymous):

Yes,i just got it thank you.